a trinangle has a 121 angle and a 5 degree angle . what is the measure of the 3rd angle?
Answer by xcentaur(357)

angles of a triangle always sum up tp 180 degrees
so let the remaining angle be x

then we get
121+5+x=180
126+x=180
x=180-126=54 degrees


Hope this helps,
Prabhat